The perimeter of a triangle is 69 cm. If the three sides of the trianglehave lengths of (3x+2), (2x-1), (4x+5), find the length of the shortest side.
Equation________________
Solution__________________

Answer by checkley71(8403)

3x+2+2x-1+4x+5=69
9x=69-6
9x=63
x=63/9
x=7
thus the shortest side is:
2x-1=2*7-1=14-1=13 shortest
3*7+2=21+2=23
4*7+5=28+5=33